The cost C (in dollars) of manufacturing x radios at Abdul's Stereo World is given by the function C(x)=0.6x^2-420x+81,763. What is the minimum cost of manufacturing radios?
Do not round your answer

Answer by lwsshak3(11628)

The cost C (in dollars) of manufacturing x radios at Abdul's Stereo World is given by the function C(x)=0.6x^2-420x+81,763. What is the minimum cost of manufacturing radios?
Do not round your answer
***
C(x)=0.6x^2-420x+81,763.
complete the square:
C(x)=0.6(x^2-(420/0.6)x)+81,763.
C(x)=0.6(x^2-700x+350^2)-73500+81,763.
C(x)=0.6(x-350)^2+8263
This is an equation of a parabola that opens upwards with vertex at (350,8263)
What is the minimum cost of manufacturing radios? 8263 at a production of 350 radios